OpenStudy (anonymous):

factor 8x6 – 56x3 + 98.

OpenStudy (anonymous):

what are the powers?/////////

OpenStudy (anonymous):

8x^6 -56x^3+98

OpenStudy (anonymous):

@nitz you their

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

can you help me with this problem @nitz

OpenStudy (amistre64):

its a quadratic in disguise

OpenStudy (amistre64):

say; u = x^3 then u^2 = x^6

OpenStudy (amistre64):

8u2 – 56u + 98 factor the u parts; then resub back in the xs

OpenStudy (anonymous):

i dont get it :(

OpenStudy (amistre64):

i assumed the ^3 and ^6 were throwing you off; but other than that, its just a quadratic to factor as usual

OpenStudy (anonymous):

it is all throwing me off i dont understand this question at all

OpenStudy (amistre64):

to factor is to pull it apart into its multipication parts

OpenStudy (amistre64):

like: 15 factors int 3 and 5

OpenStudy (amistre64):

its a good bet that an 8 factors out of this to reduce the numbers

OpenStudy (amistre64):

8u2 – 56u + 98 8 (u2 -7u +1...) 12 ----- 8)98 8 ---- 18 16 ---- 2 but maybe not

OpenStudy (amistre64):

you got it typed up correctly?

OpenStudy (anonymous):

this is the question Explain, in complete sentences, how you would factor 8x6 – 56x3 + 98. Provide the factors.

OpenStudy (amistre64):

see if 8*98 has factors that add to 56

OpenStudy (anonymous):

what when multiplied = 49 but when added equals -28 ? that is where i am stuck

OpenStudy (amistre64):

8*98 = 720 64 ----- 784 factors of 784 are: 1 784 2 392 4 98 8 49 7 112 14 56 28 28 well, that doesnt look promising

OpenStudy (amistre64):

lets try to pull it apart by completing the square 8u2 – 56u + 98 8(u2 -7u) + 98 8(u2 -7u + 3.5^2) + 98 - 8(3.5^2) 8(u-3.5)^2

OpenStudy (amistre64):

and since u=x^3 8(x^3-3.5)^2 thats about the best I can make of it :/

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=8%28x%5E3-3.5%29%5E2 the wolf likes it